OpenStudy (anonymous):

negative 4 square root of x plus 9 equals 20

OpenStudy (anonymous):

u want value of x?

OpenStudy (anonymous):

\[-4\sqrt{x+9}=20\]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

first square both sides to get rid of square root

OpenStudy (anonymous):

x =16

OpenStudy (anonymous):

Can you show me the steps please? I would like to learn how to do this on my own.

OpenStudy (anonymous):

\[-4\sqrt{x+9 }=20\] \[\sqrt{x+9}=20\div-4\] 9shifting -4 from left to right Now squaring on both sides \[(\sqrt{x+9})^{2}=(-5)^{2}\] x+9=25 x=25-9 x=16

OpenStudy (anonymous):

is it clear??

OpenStudy (anonymous):

Yes, I understand now, thank you. :)

OpenStudy (anonymous):

u welcum